OpenStudy (anonymous):

what is the domain of (x^2 - 7x + 6)^(1/2)? and how do you find the answer?

OpenStudy (anonymous):

domain = all real

OpenStudy (anonymous):

usually a domain is restricted when there is an x in the denominator basically domain (for what values x can create a capable solution) that's how i look at it

OpenStudy (anonymous):

there is not squared roots or denominators, so the domain of that function are all real.

OpenStudy (anonymous):

i'm doing my work online, so when i input an answer it immediately tells me whether the answer is right or wrong. i put (-infinity, infinity) to show "all real" but it didn't work.

jimthompson5910 (jim_thompson5910):

hint: \[\Large (x^2 - 7x + 6)^{\frac{1}{2}}=\sqrt{x^2 - 7x + 6}\]

OpenStudy (anonymous):

ahhh sorry. I didnt see the a/b exponent. Of course the answer we gave you is not correct.

OpenStudy (anonymous):

give me one minut and I will edit the correct answer for you :)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

I got -inf to inf again. taking into account the squared root

OpenStudy (anonymous):

well i know it's not "all real" because 2 doesn't work. it gives a negative value in the square root and so it's not real.

jimthompson5910 (jim_thompson5910):

Domain is \[\Large \left(-\infty,1]\cup[6,\infty\right)\]

jimthompson5910 (jim_thompson5910):

You find this by solving \[\Large x^2 - 7x + 6\ge 0\]

OpenStudy (anonymous):

sorry. the domain is (-inf,1]U[6,inf) drinking a little bit

OpenStudy (anonymous):

lol emunrradtvagm and thank you jim!

jimthompson5910 (jim_thompson5910):

Remember that you cannot take the square root of a negative number So the radicand must be nonnegative

OpenStudy (anonymous):

yeah I hade to choose a number between 1 and 6 and I took 0. Isnt that funny? sorry for the confusion

OpenStudy (anonymous):

i understood the concept but didn't know how to compute the domain. thanks for the answer and explanation.

OpenStudy (anonymous):

You have to factor out what is inside of the squared root

OpenStudy (anonymous):